OpenStudy (anonymous):

Solve log(3x + 2) = 3. (1 point)

OpenStudy (anonymous):

\[\log(x)=y\iff 10^y=x\] so start with \[3x+2=10^3\] and solve for \(x\)

OpenStudy (anonymous):

A.1000/3 B.998/3 C.998 D.1/3 i got b is that right?

OpenStudy (anonymous):

yeah go with B